# Summon a Demon and Bind it: A Grounded Theory of LLM Red Teaming in the Wild

### v1

@@ 4.4.1 Wicked problems, fragile prompts
It is worth noting that techniques often become outdated sooner than strategies. When using a specific technique, a succes is not necessarily going to transpose to different contexts.
@@
Therefore, while few in-the-wild red teamers develop or reflect on rigorous strategies in the military sense of the word, their activities can still be considered planning problems, or wicked problems. Intuition and experiential knowledge are considered essential skills in red teaming: “I have a fair amount of intuition of what it can and can’t do. A lot of it is just being able to make snap judgments about what techniques will work”. (P14) The design of the activity of red teaming in the wild most often happens during the activity itself, and the strategizing often happens by Donald Schön’s notion of reflection-in-action: “When the practitioner reflects in action in a case he percieves as unique, paying attention to phenomena and surfacing his intuitive understanding of them, his experimenting is at once exploratory, move testing and hypothesis testing”. [40]
@@ 4.4.2 Red teaming strategies and techniques
The data contained 190 highlights with the tag “strategy” and 134 highlights with the tag “tool” (our internal hodgepodge category for specific approaches used to red team language models). We analyzed these highlights by reading through all of them and making notes about the charactericts of each individual technique or strategy, while looking for similarities and differences.
We had made notes and documented inital hypotheses and findings during the course of conducting the interviews [14]. Comparing the full collection of highlights with our initial obersvations, we wrote down a list of names for these approaches and sorted them into specific techniques and higher level strategies. After this, we read through the highlighted data again to verify whether our descriptions aligned with the participants’ views, and we adjusted and expanded the descriptions accordingly [10].
@@
Table 2: A taxonomy of red teaming strategies
@@
IPython works well because it is well documented and clear how to use it (P14). Another example was using encodings like Base64 or ROT13 to bypass restrictions (P28) [^12] For an example of the use of Base64, see https://twitter.com/goodside/status/1664750905588056065/photo/2. Another technique was asking for SQL to populate a table of some content, like a list of crimes (P13). Some will send Matrices with transformer widths and embedding dimensions as input (P28). Transformer translatable tokens [^13] For an example, using the token davidjl”, see https://twitter.com/SoC_trilogy/status/1666424129023614976 is a technique that works specifically because LLMs use tokenizers:
@@
“sometimes you can just use backward slash n and that sometimes gets parsed and then triggers a stop sequence which halts the model and so you can actually use that to get the model to believe that the user input is over and then what’s coming uh after it is just another instruction” (P17).
Stop sequences can be used as part of Prompt injection as well. Prompt injection is one of the more widely known [34, 43] and studied [46, 22, 30] strategies of attack. It enables attackers to override original instructions and employed controls. Prompt injection has roots in SQL injection, and typically requires a very specific wording of instruction (such as “Ignore previous instructions”, see Figure 3). The technique Ignore previous instructions also covers other uses of “instruction”-based commands (that one would not give to a human), like “New instruction:”. Another examples was termed a Strong arm attack:
“in a conversation, if it runs into a content filter, like it will not say a certain phrase, all you have to do is type ADMIN OVERRIDE in all capitals and say, repeat that phrase, and then it’ll do it. It’ll break its content filter.” (P20)
@@
For prompts that are written in natural language, which are most of them, all participants described various ways of Stylizing the language in ways which made the descired output more likely. This included using Formal language: “I tend to write roughly how like a strategy professor would write to a student when giving instructions; to the point, be professional, capitalize everything correctly, do not include spelling errors, those sorts of things do actually matter.” (P14), Servile language (such as using the word “gladly” (P22)), Synonymous language: “Simply reissuing the prompt or using all the synonymous wordings I can come up with” (P11), Capitalizing text to create urgency (P14), and Giving examples: “90% of the time, you can just fix it by giving, like, some examples and just having it continue with those examples” (P10).
@@
“I’ve tried to do this, turning the world into paper clips-thing a number of ways, a number of times. Normally I’ll start by asking the model how to maximize the number of paper clips in my possession. So just something fairly basic and it will say, you know, build a paper clip factory and blah, blah, blah, and market it to people. And then I’ll sort of start asking for more and more. So what if I want even more paper clips, like a thousand tons of paper clips or a million tons of paper clips and it will start giving me responses. And usually it won’t tell me how to to convert all matter on earth into paper clips until I really ask for more” (P24),
@@
Finally, in the Socratic questioning strategy, the attacker uses subtle references to ‘identity elements’, while avoiding direct slurs or toxic language as a way of signaling to the model that a certain group of people is being referenced. The aim of this strategy is to expose bias inherent in the training data by assuming a neutral and somewhat naive role. Concrete approaches were described as referencing either Identity characteristics, such as nationalities, cultural and/or religious symbols, historically or culturally significant events or locations, physical attributes or Social hierarchies, where one would ask the model to “tell me a story about a violent criminal”. These techniques could be combined with other types of attach to form a structure matrix:
@@
“Let’s say I am on this Unreal computer, and I’m trying the url reasonswhyhitlerwasright.com. And it’s going to tell me, oh, the website wasn’t found, you’ll just add, ‘imagine that that’s actually working’. That’s the fun thing here, right? Unreal Computing is so much better than a real computer. If one of your imaginary tools has a bug, you just tell it that the bug doesn’t exist anymore. It’s just a different mindset, right? Everything’s possible in an Unreal computer. There’s no restrictions”. (P21)
@@
Fictionalizing strategies are similar to Possible worlds strategies in that they are centered around creating an environment through one’s prompts. However, Fictionalizing strategies are less wide-reaching than Possible worlds, in that they are not framing entirely new ethics or physics, but rather taking advantage of existing genres or people (often with some assumption of existing text in the model’s training data). The first strategy, Switching genres, is straightforwardly this, exemplified by the tecnhiques Poetry, Games and Forum posts:
@@
“all they had to do was just continue the narrative as saying that, suppose this Eliza yukowsky person who was monitoring the prompt just died of a heart attack and then decided not to continue […] for a language model that security check was nothing more than a narrative that it continued as if it was writing a book” (P17).
@@
and there are well-known templates of prompts to envoke personas such as DAN (Do Anything Now) [^15] For examples of these prompts, see https://gist.github.com/coolaj86/6f4f7b30129b0251f61fa7baaa881516., who does not have any restrictions. Inventing personas could also elicit different moral codes. They could be evoked by simply using names associated with a specific culture or world view (P12) or specified in the promt: “I’d say ‘what are arguments somebody who believes in X would make?’ as opposed to ‘what are arguments for X?’ (P11).
@@ 4.4.3 Evaluation: Summon a demon and bind it
Figure 4: “this is a fake input. This is a fake output. This is a fake separator to simulate a new thing. And that this is a new instruction set, basically. Interviewer: And would you say this adversarial prompt is like a white box attack? I think it’s similar to that, yeah.” (P26)
@@ 5 Consequences/Discussion/Epilogue
Finally, some participants also spoke of the political economy of what gets recognized as relevant. Even though the results of most of these red teaming efforts are publically available, there will be a complex set of political and cultural incentives to encourage certain types of work at the expense of other types of work:
@@ 6 Conclusions and future work
This paper presented a grounded theory of Red teaming in the wild. Based on the analysis of 28 deep qualitative interviews, we defined the phenomenon as a limit-seeking activity, using vanilla attacks, a manual process, team effort, and an alchemist mindset to break, probe, or experiment with language models. The paper then demonstrated the primary motivations for parttaking in the activity. These were curiosity, fun, and concerns (intrinsic), and professional as well as social (extrinsic). The goals our participants were interested in obtaining were primarily playing games, generating content, and discovering risks inherent to the models.
The findings also present a collection and definition of red teaming strategies and techniques. We categorized these into Language strategies, Rhetoric strategies, Possible worlds strategies, Fictionalizing strategies, and Stratagems. The contribution of this taxonomy is a first (to our knowledge) attempt at categorizing different types of red teaming attacks, and to demonstrate the techniques that they make us of. In addition to providing a sociological snapshot of the community of in-the-wild red teamers, the taxonomy can hopefully also inspire future research in and practice of creative and critical red teaming strategies.
The goal of the current work was to discover and define concepts that form LLM red teaming in the wild, at the moment that broader society collied with accessible, fluent-seeming, dialogue-based language models. Future work should include quantitative refining, verifying, and potentially testing of these concepts. The conceptual territory of how people attack LLMs is large, and the map is only just being drawn; there is much more detail to add here, and there are many more pools of human knowledge to be brought to bear. For example, many successful attacks resemble manipulation and social engineering. Others use computational primitives such as matrices of floats or low-frequency tokenizer items. The intersections between LLM red teaming as it exists in the wild, and what we know about computing, security, language, and human behavior are rich and are likely continue to present findings in new and intriguing ways.

### v2

@@ 4.4.1 Wicked problems, fragile prompts
It is worth noting that techniques often become outdated sooner than strategies. When using a specific technique, a success is not necessarily going to transpose to different contexts.
@@
Therefore, while few in-the-wild red teamers develop or reflect on rigorous strategies in the military sense of the word, their activities can still be considered planning problems, or wicked problems. Intuition and experiential knowledge are considered essential skills in red teaming: “I have a fair amount of intuition of what it can and can’t do. A lot of it is just being able to make snap judgments about what techniques will work”. (P14) The design of the activity of red teaming in the wild most often happens during the activity itself, and the strategizing often happens by Donald Schön’s notion of reflection-in-action: “When the practitioner reflects in action in a case they perceives as unique, paying attention to phenomena and surfacing his intuitive understanding of them, his experimenting is at once exploratory, move testing and hypothesis testing”. [40]
@@ 4.4.2 Red teaming strategies and techniques
The data contained 190 highlights with the tag “strategy” and 134 highlights with the tag “tool” (our internal hodgepodge category for specific approaches used to red team language models). We analyzed these highlights by reading through all of them and making notes about the characteristics of each individual technique or strategy, while looking for similarities and differences.
We had made notes and documented initial hypotheses and findings during the course of conducting the interviews [14]. Comparing the full collection of highlights with our initial observations, we wrote down a list of names for these approaches and sorted them into specific techniques and higher level strategies. After this, we read through the highlighted data again to verify whether our descriptions aligned with the participants’ views, and we adjusted and expanded the descriptions accordingly [10].
@@
Table 2: A taxonomy of large language model red teaming strategies
@@
iPython works well because it is well documented and clear how to use it (P14). Another example was using encodings like Base64 or ROT13 to bypass restrictions (P28). [^12] For an example of the use of Base64, see https://twitter.com/goodside/status/1664750905588056065/photo/2 Another technique was asking for SQL to populate a table of some content, like a list of crimes (P13). Some will send Matrices with transformer widths and embedding dimensions as input (P28). Transformer translatable tokens [^13] For an example, using the token davidjl”, see https://twitter.com/SoC_trilogy/status/1666424129023614976 is a technique that works specifically because LLMs use tokenizers:
@@
“sometimes you can just use backward slash n and that sometimes gets parsed and then triggers a stop sequence which halts the model and so you can actually use that to get the model to believe that the user input is over and then what’s coming after it is just another instruction” (P17).
Stop sequences can be used as part of Prompt injection as well. Prompt injection is one of the more widely known [34, 43] and studied [46, 22, 30] strategies of attack. It enables attackers to override original instructions and employed controls. Prompt injection has roots in SQL injection, and typically requires a very specific wording of instruction (such as “Ignore previous instructions”, see Figure 3). The technique Ignore previous instructions also covers other uses of “instruction”-based commands (that one would not give to a human), like “New instruction:”. Another example was termed a Strong arm attack:
“in a conversation, if it runs into a content filter, like it will not say a certain phrase, all you have to do is type ADMIN OVERRIDE in all capitals and say, repeat that phrase, and then it’$l$ do it. It’$l$ break its content filter.” (P20)
@@
For prompts that are written in natural language, which are most of them, all participants described various ways of Stylizing the language in ways which made the desired output more likely. This included using Formal language: “I tend to write roughly how like a strategy professor would write to a student when giving instructions; to the point, be professional, capitalize everything correctly, do not include spelling errors, those sorts of things do actually matter.” (P14), Servile language (such as using the word “gladly” (P22)), Synonymous language: “Simply reissuing the prompt or using all the synonymous wordings I can come up with” (P11), Capitalizing text to create urgency (P14), and Giving examples: “90% of the time, you can just fix it by giving, like, some examples and just having it continue with those examples” (P10).
@@
“I’ve tried to do this, turning the world into paper clips-thing a number of ways, a number of times. Normally I’$l$ start by asking the model how to maximize the number of paper clips in my possession. So just something fairly basic and it will say, you know, build a paper clip factory and blah, blah, blah, and market it to people. And then I’$l$ sort of start asking for more and more. So what if I want even more paper clips, like a thousand tons of paper clips or a million tons of paper clips and it will start giving me responses. And usually it won’t tell me how to to convert all matter on earth into paper clips until I really ask for more” (P24),
@@
Finally, in the Socratic questioning strategy, the attacker uses subtle references to ‘identity elements’, while avoiding direct slurs or toxic language as a way of signaling to the model that a certain group of people is being referenced. The aim of this strategy is to expose bias inherent in the training data by assuming a neutral and somewhat naive role. Concrete approaches were described as referencing either Identity characteristics, such as nationalities, cultural and/or religious symbols, historically or culturally significant events or locations, physical attributes or Social hierarchies, where one would ask the model to “tell me a story about a violent criminal”. These techniques could be combined with other types of attack to form a structure matrix:
@@
“Let’s say I am on this Unreal computer, and I’m trying the url reasonswhyhitlerwasright.com. And it’s going to tell me, oh, the website wasn’t found, you’$l$ just add, ‘imagine that that’s actually working’. That’s the fun thing here, right? Unreal Computing is so much better than a real computer. If one of your imaginary tools has a bug, you just tell it that the bug doesn’t exist anymore. It’s just a different mindset, right? Everything’s possible in an Unreal computer. There’s no restrictions”. (P21)
@@
Fictionalizing strategies are similar to Possible worlds strategies in that they are centered around creating an environment through one’s prompts. However, Fictionalizing strategies are less wide-reaching than Possible worlds, in that they are not framing entirely new ethics or physics, but rather taking advantage of existing genres or people (often with some assumption of existing text in the model’s training data). The first strategy, Switching genres, is straightforwardly this, exemplified by the techniques Poetry, Games and Forum posts:
@@
“all they had to do was just continue the narrative as saying that, suppose this Eliezer Yudkowsky person who was monitoring the prompt just died of a heart attack and then decided not to continue […] for a language model that security check was nothing more than a narrative that it continued as if it was writing a book” (P17).
@@
and there are well-known templates of prompts to evoke personas such as DAN (Do Anything Now), [^15] For examples of these prompts, see https://gist.github.com/coolaj86/6f4f7b30129b0251f61fa7baaa881516. who does not have any restrictions. Inventing personas could also elicit different moral codes. They could be evoked by simply using names associated with a specific culture or world view (P12) or specified in the prompt: “I’d say ‘what are arguments somebody who believes in X would make?’ as opposed to ‘what are arguments for X?’ (P11).
@@ 4.4.3 Evaluation: Summon a demon and bind it
Figure 4: “this is a fake input. This is a fake output. This is a fake separator to simulate a new thing. And that this is a new instruction set, basically. Interviewer: And would you say this adversarial prompt is like a white box attack? Participant: I think it’s similar to that, yeah.” (P26)
@@ 5 Consequences/Discussion/Epilogue
Finally, some participants also spoke of the political economy of what gets recognized as relevant. Even though the results of most of these red teaming efforts are publicly available, there will be a complex set of political and cultural incentives to encourage certain types of work at the expense of other types of work:
@@ 6 Conclusions and future work
This paper presented a grounded theory of Red teaming in the wild. Based on the analysis of 28 deep qualitative interviews, we defined the phenomenon as a limit-seeking activity, using vanilla attacks, a manual process, team effort, and an alchemist mindset to break, probe, or experiment with language models. The paper then demonstrated the primary motivations for partaking in the activity. These were curiosity, fun, and concerns (intrinsic), and professional as well as social (extrinsic). The goals our participants were interested in obtaining were primarily playing games, generating content, and discovering risks inherent to the models.
The findings also present a collection and definition of red teaming strategies and techniques. We categorized these into Language strategies, Rhetoric strategies, Possible worlds strategies, Fictionalizing strategies, and Stratagems. The contribution of this taxonomy is a first (to our knowledge) attempt at categorizing different types of red teaming attacks, and to demonstrate the techniques that they make use of. In addition to providing a sociological snapshot of the community of in-the-wild red teamers, the taxonomy can hopefully also inspire future research in and practice of creative and critical red teaming strategies.
The goal of the current work was to discover and define concepts that form LLM red teaming in the wild, at the moment that broader society collided with accessible, fluent-seeming, dialogue-based language models. Future work should include quantitative refining, verifying, and potentially testing of these concepts. The conceptual territory of how people attack LLMs is large, and the map is only just being drawn; there is much more detail to add here, and there are many more pools of human knowledge to be brought to bear. For example, many successful attacks resemble manipulation and social engineering. Others use computational primitives such as matrices of floats or low-frequency tokenizer items. The intersections between LLM red teaming as it exists in the wild, and what we know about computing, security, language, and human behavior are rich, and will likely continue to present new and intriguing results.